The ABM club wants to rent a bus for the yearly

educational trip. The rent of the bus is P10,000,

which will be divided among the members that agree

to join. At this moment, 8 students have already

signed up.

a. at this moment, what is the cost for each member?

b. by making a table, show what happen to the cost

per member as additional members agrees to join.

c. Write a function for the cost per member C(m)

related to the number of additional members m that

agree to join.


1
Expert's answer
2021-10-04T16:39:13-0400

(a)

"\dfrac{10000}{8}=1250"

The cost for each member is P1250.

(b) Let "m=" the number of additional members that agree to join.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 m & C(m)\\ \hline
 0 & 1250\\
 1 & 10000/9\approx1111.11\\
 2 & 1000\\
 3 & 10000/11\approx909.09\\
 4 & 10000/12\approx833.33\\
 5 & 10000/13\approx769.23\\
 6 & 10000/14\approx714.29\\
 7 & 10000/15\approx666.67\\
 8 & 625\\
 9 & 10000/17\approx588.24\\
 10 & 10000/18\approx555.56\\
 11 & 10000/19\approx526.32\\
 12 & 500\\
\end{array}"

The cost per member "C(m)" decreases from P1250 if the number of additional members "m" increases from "0."

(c)


"C(m)=\dfrac{10000}{8+m}, m\geq 0"